# The Christoffel problem for the disk area measure

### v2

@@ 2.2. Disintegration of measures
Next, we want to show that under the given assumptions on the pushforward of the measure $\mu$ in Theorem˜A, there exists a disintegration ˜1.5. To this end, we recall the disintegration theorem from measure theory (see, e.g., [Ambrosio2008]*Theorem 5.3.1).
@@
First, we let $\nu=\pi_{\ast}\mu$ and denote by $(\tilde{\mu}_{E})_{E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}$ the disintegration of $\mu$ according to ˜2.4. By our assumption on $\mu$, we have $\nu(dE)=\rho(E)dE$ with some density $\rho\in C(\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n}))$. Due to the continuity of $\rho$, we may now define $\mu_{E}:=\rho(E)\tilde{\mu}_{E}$ for $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$. It is then easy to see that the family $(\mu_{E})$ fulfills the assertion. ∎
2.3. Proof of Theorem˜A
Now we are ready to prove the main result of this article, Theorem˜A. In the proof below, we readily identify $\mu$ with its restriction to $\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$, and denote by $(\mu_{E})$ its disintegration according to ˜2.5.
Proof of Theorem˜A.
To show that the conditions ˜(i) to ˜(iii) are necessary, suppose that there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $\mu=S_{1}(K,\mathbb{D},{}\cdot{})$. We identify $\mu$ with its restriction to $\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$, take some function $\tilde{f}\in C(\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n}))$ and let $f:=\tilde{f}\circ\pi$, that is, $f(u)=\tilde{f}(e_{n}\vee u)$. Then the Cauchy–Kubota type formula ˜2.1, combined with ˜2.2, gives
@@ 2.3. Proof of Theorem˜A
and thus, $\pi_{\ast}\mu=\kappa_{n-1}V_{1}(K|E)dE$. Since the expression $V_{1}(K|E)$ is continuous in $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$, the push-forward measure $\pi_{\ast}\mu$ is absolutely continuous with a continuous density, showing ˜(i). Moreover, due to ˜2.1 and the uniqueness in ˜2.5, its disintegration $(\mu_{E})_{E}$ is given by $\mu_{E}=\frac{\kappa_{n-1}}{2}S_{1}^{E}(K|E,{}\cdot{})$. Hence, for a.e. $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$, the measure $\mu_{E}$ (as a measure on $\mathbb{S}^{1}(E)$) is centered, showing ˜(ii). To verify the final condition ˜(iii), we let $h$ be the support function of $K$ and combine ˜1.3 and ˜1.1 in dimension $2$ to obtain
$h(u)-\frac{1}{\pi}\int_{\mathbb{S}^{1}(e_{n}\vee u)}h(v)\langle u,v\rangle\,dv=h_{K|(e_{n}\vee u)-s(K|(e_{n}\vee u))}(u)$
$\qquad=\frac{2}{\pi\kappa_{n-1}}\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\langle u,v\rangle^{2}}(\pi-\arccos\langle u,v\rangle)\,\mu_{e_{n}\vee u}(dv).$
Conversely, suppose that $\mu$ is a non-negative finite Borel measure on $\mathbb{S}^{n-1}$ satisfying the conditions ˜(i) to ˜(iii), with $(\mu_{E})_{E}$ denoting its disintegration according to ˜2.5. Moreover, take $K\in\mathcal{K}(\mathbb{R}^{n})$ to be the convex body that has $h$ as its support function. For a.e. subspace $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$, the measure $\mu_{E}$ is centered, so by ˜1.1, there exists a convex body $K_{E}\in\mathcal{K}(E)$ such that $\mu_{E}=S_{1}^{E}(K_{E},{}\cdot{})$. Combining ˜1.3 and ˜1.2, we obtain that
$h_{K|E-s(K|E)}(u)=\int_{\mathbb{S}^{1}(E)}\sqrt{1-\langle u,v\rangle^{2}}(\pi-\arccos\langle u,v\rangle)\,\mu_{E}(dv)=\pi h_{K_{E}-s(K_{E})}(u)$
for every $u\in\mathbb{S}^{1}(E)$. This shows that $K|E-s(K|E)=\pi(K_{E}-s(K_{E}))$ and thus, $\mu_{E}=S_{1}^{E}(K_{E},{}\cdot{})=\frac{1}{\pi}S_{1}^{E}(K|E,{}\cdot{})$. As $\mu_{E}$ has no mass at $\pm e_{n}$, so has $S_{i}^{E}(K|E,{{}\cdot{}})$. Therefore, combining ˜2.1 and ˜2.6, we have that for every bounded Borel function $f:\mathbb{S}^{n-1}\to\mathbb{R}$,
@@
For the uniqueness part of the statement, take two convex bodies $K,L\in\mathcal{K}(\mathbb{R}^{n})$ such that $S_{1}(K,\mathbb{D},{}\cdot{})=S_{1}(L,\mathbb{D};{}\cdot{})$, and also, $S_{1}(K,\mathbb{D},\{\pm e_{n}\})=0$. By integrating the respective support function of $K$ and $L$ against these mixed area measures, by ˜2.2, we have that $V(K,K,\mathbb{D}^{[n-2]})=V(K,L,\mathbb{D}^{[n-2]})=V(L,L,\mathbb{D}^{[n-2]})$.
Thus, equality is attained in Minkowski’s quadratic inequality ˜2.5 for $C=\mathbb{D}$, so according to ˜2.3, the bodies $K+F(L,e_{n})$ and $L+F(K,e_{n})$ are equal. Noting, however, that due to Lemma˜2.1 and identities ˜2.3 and ˜2.4,
@@
Under the addition that $\mu$ is even, which corresponds to $K$ being centrally symmetric, the classification simplifies considerably. This is the content of the theorem below, which we obtain from an easy modification of the proof of Theorem˜A.
@@
(2.7) $h(u)=\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\langle u,v\rangle^{2}}\,\mu_{e_{n}\vee u}(dv),$
@@
Suppose first that there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $\mu=S_{1}(K,\mathbb{D},{}\cdot{})$. By the evenness of $\mu$ and compatibility of mixed area measures with linear isometries, $\mu=S_{1}(-K,\mathbb{D},{}\cdot{})$. By the uniqueness statement in Theorem˜A, there exists some $x\in\mathbb{R}^{n}$ such that K=xKK=x-K, that is, $K$ is symmetric about $x$.
Due to Theorem˜A, the push-forward measure $\pi_{\ast}\mu$ is absolutely continuous with a continuous density, and from the proof of Theorem˜A, the disintegration of $\mu$ is necessarily given by $\mu_{E}=\frac{\kappa_{n-1}}{2}S_{1}^{E}(K|E,{}\cdot{})$ for a.e. $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$. In particular, $\mu_{E}$ (as a measure on $\mathbb{S}^{1}(E)$) is even for a.e. $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$. Moreover, from the proof of Theorem˜A, the support function $h$ of a dilated copy of $K$ satisfies identity ˜1.6. Due to the evenness of $h$ and $\mu_{E}$, this identity then simplifies to ˜2.7.
@@
Consequently, the right hand side of ˜2.7 is an even function of $u$, and thus, so is the left hand side. Since identity ˜2.7 is simply the even component of identity ˜1.6, we deduce that $h$ also satisfies ˜1.6. In conclusion, $\mu$ meets the requirements of Theorem˜A, and thus, there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $\mu=S_{1}(K,\mathbb{D},{}\cdot{})$. ∎
@@
$\int_{\mathbb{S}^{n-1}}q(u)\,du=\frac{1}{2}\int_{\mathbb{S}^{n-2}(e_{n}^{\perp})}\int_{\mathbb{S}^{1}(e_{n}\vee u)}q(v)\lvert\langle u,v\rangle\rvert^{n-2}\,dv\,du.$
From this formula, it is immediate that $\pi_{\ast}\mu$ is absolutely continuous with a continuous density given by $\frac{d\pi_{\ast}\mu}{dE}=\frac{\omega_{n-1}}{2}\int_{\mathbb{S}^{1}(E)}q(v)\lVert v|E\rVert^{n-2}\,dv$. Moreover, the disintegration of $\mu$ in the sense of ˜2.6 is given by
$\mu_{E}(dv)=\frac{\omega_{n-1}}{2}q(v)\lVert v|E\rVert^{n-2}\,dv.$
Hence, as a special case of Theorem˜A, we obtain the following.
@@
Let $q$ be a non-negative $L^{1}(\mathbb{S}^{n-1})$ function. Then there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $q(u)du=S_{1}(K,\mathbb{D},du)$ if and only if $\int_{\mathbb{S}^{1}(e_{n}\vee u)}q(u)\lvert\langle u,v\rangle\rvert^{n-2}v\,dv=0$ for almost every $u\in\mathbb{S}^{n-2}(e_{n}^{\perp})$, and there exists a support function $h$ such that
$h(u)-\frac{1}{\pi}\int_{\mathbb{S}^{1}(e_{n}\vee u)}h(v)\langle u,v\rangle\,dv=\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\langle u,v\rangle^{2}}(\pi-\arccos\langle u,v\rangle)\lvert\langle u,v\rangle\rvert^{n-2}q(v)\,dv$
@@
Under the additional assumption of evenness, ˜2.6 yields the following.
@@
$h(u)=\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\langle u,v\rangle^{2}}\lvert\langle u,v\rangle\rvert^{n-2}q(v)\,dv,$

### v3

@@ 2.2. Disintegration of measures
Next, we want to show that under the given assumptions on the pushforward of the measure $\mu$ in ˜A, there exists a disintegration ˜1.5. To this end, we recall the disintegration theorem from measure theory (see, e.g., [Ambrosio2008]*Theorem 5.3.1).
@@
First, we let $\nu=\pi_{\ast}\mu$ and denote by $(\tilde{\mu}_{E})_{E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})}$ the disintegration of $\mu$ according to Theorem˜2.4. By our assumption on $\mu$, we have $\nu(dE)=\rho(E)dE$ with some density $\rho\in C(\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n}))$. Due to the continuity of $\rho$, we may now define $\mu_{E}:=\rho(E)\tilde{\mu}_{E}$ for $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$. It is then easy to see that the family $(\mu_{E})$ fulfills the assertion. ∎
2.3. Proof of ˜A
Now we are ready to prove the main result of this article, ˜A. In the proof below, we readily identify $\mu$ with its restriction to $\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$, and denote by $(\mu_{E})$ its disintegration according to ˜2.5.
Proof of ˜A.
To show that the conditions ˜(i) to ˜(iii) are necessary, suppose that there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $\mu=S_{1}(K,\mathbb{D},{}\cdot{})$. We identify $\mu$ with its restriction to $\mathbb{S}^{n-1}\setminus\{\pm e_{n}\}$, take some function $\tilde{f}\in C(\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n}))$ and let $f:=\tilde{f}\circ\pi$, that is, $f(u)=\tilde{f}(e_{n}\vee u)$. Then the Cauchy–Kubota type formula ˜2.1, combined with ˜2.3, gives
@@ 2.3. Proof of ˜A
and thus, $\pi_{\ast}\mu=\kappa_{n-1}V_{1}(K|E)dE$. Since the expression $V_{1}(K|E)$ is continuous in $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$, the push-forward measure $\pi_{\ast}\mu$ is absolutely continuous with a continuous density, showing ˜(i). Moreover, due to ˜2.1 and the uniqueness in ˜2.5, its disintegration $(\mu_{E})_{E}$ is given by $\mu_{E}=\frac{\kappa_{n-1}}{2}S_{1}^{E}(K|E,{}\cdot{})$. Hence, for a.e. $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$, the measure $\mu_{E}$ (as a measure on $\mathbb{S}^{1}(E)$) is centered, showing ˜(ii). To verify the final condition ˜(iii), we let $h$ be the support function of $K$ and combine ˜2.2 and Theorem˜1.1 in dimension $2$ to obtain
$h(u)-\frac{1}{\pi}\int_{\mathbb{S}^{1}(e_{n}\vee u)}h(v)\left\langle u,v\right\rangle\,dv=h_{K|(e_{n}\vee u)-s(K|(e_{n}\vee u))}(u)$
$\qquad=\frac{2}{\pi\kappa_{n-1}}\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\left\langle u,v\right\rangle^{2}}(\pi-\arccos\left\langle u,v\right\rangle)\,\mu_{e_{n}\vee u}(dv).$
Conversely, suppose that $\mu$ is a non-negative finite Borel measure on $\mathbb{S}^{n-1}$ satisfying the conditions ˜(i) to ˜(iii), with $(\mu_{E})_{E}$ denoting its disintegration according to ˜2.5. Moreover, take $K\in\mathcal{K}(\mathbb{R}^{n})$ to be the convex body that has $h$ as its support function. For a.e. subspace $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$, the measure $\mu_{E}$ is centered, so by Theorem˜1.1, there exists a convex body $K_{E}\in\mathcal{K}(E)$ such that $\mu_{E}=S_{1}^{E}(K_{E},{}\cdot{})$. Combining ˜2.2 and ˜1.3, we obtain that
$h_{K|E-s(K|E)}(u)=\int_{\mathbb{S}^{1}(E)}\sqrt{1-\left\langle u,v\right\rangle^{2}}(\pi-\arccos\left\langle u,v\right\rangle)\,\mu_{E}(dv)=\pi h_{K_{E}-s(K_{E})}(u)$
for every $u\in\mathbb{S}^{1}(E)$. This shows that $K|E-s(K|E)=\pi(K_{E}-s(K_{E}))$ and thus, $\mu_{E}=S_{1}^{E}(K_{E},{}\cdot{})=\frac{1}{\pi}S_{1}^{E}(K|E,{}\cdot{})$. As $\mu_{E}$ has no mass at $\pm e_{n}$, so has $S_{i}^{E}(K|E,{{}\cdot{}})$. Therefore, combining ˜2.1 and ˜2.7, we have that for every bounded Borel function $f:\mathbb{S}^{n-1}\to\mathbb{R}$,
@@
For the uniqueness part of the statement, take two convex bodies $K,L\in\mathcal{K}(\mathbb{R}^{n})$ such that $S_{1}(K,\mathbb{D},{}\cdot{})=S_{1}(L,\mathbb{D};{}\cdot{})$, and also, $S_{1}(K,\mathbb{D},\{\pm e_{n}\})=0$. By integrating the respective support function of $K$ and $L$ against these mixed area measures, by ˜2.3, we have that $V(K,K,\mathbb{D}^{[n-2]})=V(K,L,\mathbb{D}^{[n-2]})=V(L,L,\mathbb{D}^{[n-2]})$.
Thus, equality is attained in Minkowski’s quadratic inequality ˜2.6 for $C=\mathbb{D}$, so according to ˜2.3, the bodies $K+F(L,e_{n})$ and $L+F(K,e_{n})$ are equal. Noting, however, that due to ˜2.1 and identities ˜2.4 and ˜2.5,
@@
Under the addition that $\mu$ is even, which corresponds to $K$ being centrally symmetric, the classification simplifies considerably. This is the content of the theorem below, which we obtain from an easy modification of the proof of ˜A.
@@
(2.8) $h(u)=\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\left\langle u,v\right\rangle^{2}}\,\mu_{e_{n}\vee u}(dv),$
@@
Suppose first that there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $\mu=S_{1}(K,\mathbb{D},{}\cdot{})$. By the evenness of $\mu$ and compatibility of mixed area measures with linear isometries, $\mu=S_{1}(-K,\mathbb{D},{}\cdot{})$. By the uniqueness statement in ˜A, there exists some $x\in\mathbb{R}^{n}$ such that K=xKK=x-K, that is, $K$ is symmetric about $x$.
Due to ˜A, the push-forward measure $\pi_{\ast}\mu$ is absolutely continuous with a continuous density, and from the proof of ˜A, the disintegration of $\mu$ is necessarily given by $\mu_{E}=\frac{\kappa_{n-1}}{2}S_{1}^{E}(K|E,{}\cdot{})$ for a.e. $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$. In particular, $\mu_{E}$ (as a measure on $\mathbb{S}^{1}(E)$) is even for a.e. $E\in\mathrm{Gr}_{2}(\mathbb{R}^{n},e_{n})$. Moreover, from the proof of ˜A, the support function $h$ of a dilated copy of $K$ satisfies identity ˜1.6. Due to the evenness of $h$ and $\mu_{E}$, this identity then simplifies to ˜2.8.
@@
Consequently, the right hand side of ˜2.8 is an even function of $u$, and thus, so is the left hand side. Since identity ˜2.8 is simply the even component of identity ˜1.6, we deduce that $h$ also satisfies ˜1.6. In conclusion, $\mu$ meets the requirements of ˜A, and thus, there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $\mu=S_{1}(K,\mathbb{D},{}\cdot{})$. ∎
@@
$\int_{\mathbb{S}^{n-1}}q(u)\,du=\frac{1}{2}\int_{\mathbb{S}^{n-2}(e_{n}^{\perp})}\int_{\mathbb{S}^{1}(e_{n}\vee u)}q(v)\lvert\left\langle u,v\right\rangle\rvert^{n-2}\,dv\,du.$
From this formula, it is immediate that $\pi_{\ast}\mu$ is absolutely continuous with a continuous density given by $\frac{d\pi_{\ast}\mu}{dE}=\frac{\omega_{n-1}}{2}\int_{\mathbb{S}^{1}(E)}q(v)\lVert v|E\rVert^{n-2}\,dv$. Moreover, the disintegration of $\mu$ in the sense of ˜2.7 is given by
$\mu_{E}(dv)=\frac{\omega_{n-1}}{2}q(v)(1-\left\langle e_{n},v\right\rangle^{2})^{\frac{n-2}{2}}\,dv.$
Hence, as a special case of ˜A, we obtain the following.
@@
Let $q$ be a non-negative $L^{1}(\mathbb{S}^{n-1})$ function. Then there exists a convex body $K\in\mathcal{K}(\mathbb{R}^{n})$ such that $q(u)du=S_{1}(K,\mathbb{D},du)$ if and only if $\int_{\mathbb{S}^{1}(e_{n}\vee u)}q(v)\lvert\left\langle u,v\right\rangle\rvert^{n-2}v\,dv=0$ for almost every $u\in\mathbb{S}^{n-2}(e_{n}^{\perp})$, and there exists a support function $h$ such that
$h(u)-\frac{1}{\pi}\int_{\mathbb{S}^{1}(e_{n}\vee u)}h(v)\left\langle u,v\right\rangle\,dv$
$\qquad=\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\left\langle u,v\right\rangle^{2}}(\pi-\arccos\left\langle u,v\right\rangle)(1-\left\langle e_{n},v\right\rangle^{2})^{\frac{n-2}{2}}q(v)\,dv$
@@
Under the additional assumption of evenness, Theorem˜2.6 yields the following.
@@
$h(u)=\int_{\mathbb{S}^{1}(e_{n}\vee u)}\sqrt{1-\left\langle u,v\right\rangle^{2}}(1-\left\langle e_{n},v\right\rangle^{2})^{\frac{n-2}{2}}q(v)\,dv,$